NASA ISS On-Orbit Status 20 January 2009
January 20, 2009

All ISS systems continue to function nominally, except those noted previously or below. US Presidential Inauguration Day. [Per special crew request, President Obama’s Taking of the Oath of Office was uplinked to the ISS.]

Continuing the troubleshooting of the CGSE (Common Gas Support Equipment) in the JAXA JPM (JEM Pressurized Module), which has a blockage problem, CDR Fincke had 3:35h set aside for removing and replacing the CO2 valve with the He VU (helium valve unit) and stow the defective CO2 VU for return on 15A. [Since the SSV1 (Self Shut-off Valve 1) is one of the suspected points of the blockage, Mike installed a position holder in the He VU to keep the SSV opened permanently as FE-2 Magnus did for the CO2 VU on 12/15/08. The removal of the CO2 VU, which also removed the CGSE fireport, was determined to be acceptable since the potential risk of fire in the CGSE is extremely low and the capability to inject fire suppressant with PFE (Portable Fire Extinguisher) still exists, as does monitoring with CSA-CP (Compound Specific Analyzer-Combustion Products).]

FE-1 Lonchakov continued preparations for operating the Russian/German TEKh-20 Plasma Crystal-3 Plus (PK-3+) experiment payload, the first time for Expedition 18. [After unstowing and setting up the hardware yesterday in the Service Module (SM), leak checking of the electronics box and evacuation of the vacuum work chamber (ZB) in the SM Work Compartment (RO) with the turbopump, Lonchakov today conducted more hardware testing and calibration, uploaded new software from a USB stick, checked out the software installation and verified the readiness of the experiment. After starting the turbo pump right after wake-up and conducting additional leak checking on the ZB during the “day”, the FE-1 will deactivate the turbopump tonight at ~4:25pm EST. The resulting log file was then downloaded to laptop for downlink via BSR-TM. The experiment is performed on plasma, i.e., fine particles charged and excited by HF (high frequency) radio power inside the evacuated work chamber. Main objective is to obtain a homogeneous plasma dust cloud at various pressures and particle quantities with or without superimposition of an LF (low frequency) harmonic electrical field. The experiment is conducted in automated mode. PK-3+ has more advanced hardware and software than the previously used Russian PKE-Nefedov payload.]

After power-cycling the ARED (Advanced Resistive Exercise Device) display (which had a missing driver uplinked yesterday), FE-2 Magnus set up the G1 camcorder with two 50-ft video cables in the Lab to cover the subsequent ARED activities in Node-1. Magnus & Fincke then successfully performed the long-awaited ACO (Activation & Checkout) of the new exercise device. Later, Magnus dismantled & stowed the video equipment and returned the two camcorders to their nominal locations in the Lab and A/L (Airlock). [The exercise sessions were performed without the sensor calibration which is pending further software troubleshooting. At the conclusion of the exercise, Fincke noted a frayed cable and downlinked video and photos. Specialists will analyze the images and data and provide recommendations for the crew.]

Continuing the extended leak checking of the spare BZh Liquid Unit (#056) for the Elektron O2 generator, FE-1 Lonchakov charged the unit once again with pressurized N2 from the BPA Nitrogen Purge Unit (#23) to 1 atm (1 kg/cm2). The last test pressurization to monitor for leakage was on 12/20/08. [Objective of the monthly checkout of the BZh, which has been in stowage for about 2 years, is to check for leakage and good water passage through the feed line inside of the BZh (from ZL1 connector to the buffer tank) and to check the response of the Electronics Unit’s micro switches (signaling “Buffer Tank is Empty” & “Buffer Tank is Full”. During Elektron operation, the inert gas locked up in the BZh has the purpose to prevent dangerous O2/H2 mixing. A leaking BZh cannot be used.]

In the US A/L (Airlock), Magnus terminated the regeneration of METOX (Metal Oxide) EMU (Extravehicular Mobility Unit) CO2 removal canisters #0007 & #0011 in the “bake-out” oven and installed two more expended canisters (#0012, #0013), in preparation for the STS-119 spacewalks.

Also for the 15A spacewalks, Sandy had ~2.5 hrs scheduled in the A/L for completing Part 1 (of three) of EVA (Extravehicular Activity) Tool configuring and restowing.

Working on the CIR (Combustion Integrated Rack) in the Lab, FE-2 Magnus reversed the installed direction of the return umbilical of the ARIS MOD Temp TCS (Active Rack Isolation System Moderate Temperature Thermal Control System). She then repositioned the CIR in the rack bay by adjusting PaRIS (Passive Rack Isolation System) centering hardware (four front snubber pins/cups).

Lonchakov spent ~1.5 hrs on the TVIS treadmill for the periodic Russian PZE-MO-3 test for physical fitness evaluation, his second time, using the TVIS in unmotorized (manual control) mode and wearing the Kardiokassette KK-2000 belt with three chest electrodes. [The fitness test, controlled from the RSE-Med laptop, yields ECG (electrocardiogram) readings to the KK-2000 data storage device, later downlinked via the Regul (BSR-TM) payload telemetry channel. Before the run, the KK-2000 was synchronized with the computer date/time readings. For the ECG, the crewmembers worked out on the treadmill, first walking 3 min. up to 3.5 km/h, then running at a slow pace of 5-6 km/h for 2 min, at moderate pace of 6.5 km/h, followed by the maximum pace not exceeding 10 km/h, then walking again at gradually decreasing pace.]

Fincke initiated (later terminated) a software “ghosting” upgrade of the SAMS (Space Acceleration Measurement System) hard drive and monitored data outputs from the display screen for calldown to the ground. [While the SAMS laptop shell is functioning properly, the flight hard drives carrying the SAMS software & operating system are only marginally functional. On 1J, a Ghost Load CD and floppy disk were sent up with the intent of ghosting new hard drives to replace the marginal ones. In August ‘08, the ghosting process was started, and is currently partially complete. The reformatted hard drives are installed in the SAMS ICU (Interface Control Unit), but the ghost load which is resident on the EXPRESS-4 Rack Laptop has not yet been copied to the hard drives. When the task is completed, SAMS will be able to use the results to configure multiple sets of hard drives on orbit and will be able to provide acceleration measurement data and more continuous support to the microgravity scientific community as well as the ISS structures group.]

Sandy Magnus completed the daily flushing of the PWD (Potable Water Dispenser). [The PWD had been found, via several microbial analyses by Magnus, to have bacteria growing in the ambient water. It is suspected that this is due to the water being stagnant and not used. The crew now performs daily flushes with 100 mL of iodinated water.]

Working on the US OGS (Oxygen Generator System) in the Lab, the FE-2 removed & replaced the ACTEX filter cartridge in the Wastewater jumper. [New cartridge: #2003.]

The FE-1 completed the routine daily servicing of the SM’s SOZh system (Environment Control & Life Support System, ECLSS). [Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers, replacement of EDV-SV waste water and EDV-U urine containers and performing US condensate processing (transfer from US CWC to Russian EDV containers) if condensate is available.]

Yuri also performed the regular daily IMS (Inventory Management System) maintenance task by updating/editing the IMS standard “delta file” including stowage locations for the regular weekly automated export/import to its three databases on the ground (Houston, Moscow, Baikonur).

Continuing the current round of periodic preventive maintenance of RS (Russian Segment) ventilation systems, Mike Fincke replaced the filters of the PS1 & PS2 dust collectors in the FGB (Funktsionalnyi-Grusovoi Blok), then updated the IMS accordingly.

Working with the CDR, Yuri Lonchakov performed operational and leak tests on two solenoid (electromagnetic) valves (EK9, EK10) on the PK3 (Pneumatic Control 3) panel 9416 of the SOTR/TCS (Thermal Control System) recharge loop in the PrK (SM Transfer Compartment), in “reportage mode”.

Mike was also scheduled to perform troubleshooting in the ESA COL (Columbus Orbital Laboratory) on the DLT (Digital Line Tape) recorder of the FSL VMU (Fluid Science Laboratory/Video Management Unit).

The crew undertook another periodic relocation of the TEPC (Tissue Equivalent Proportional Counter) detector assembly, the primary radiation measurement tool in the ISS, today in Node-2 from inside the new portside CQ (Crew Quarters) to the starboard CQ. [TEPC had been moved to the portside CQ by Fincke on 12/26, after having been in the SM on Panel 327 from 11/27 and before that in the Node-2 from 11/10.]

At ~1:01pm EST, the FE-2 powered up the SM’s amateur radio equipment (Kenwood VHF transceiver with manual frequency selection, headset, & power supply) and conducted, at 1:06pm, a ham radio session with Children’s Hospital of Eastern Ontario, Ottawa, Ontario, Canada. [The Children’s Hospital of Eastern Ontario (CHEO) is an academic paediatric hospital affiliated with the University of Ottawa, with a mandate for care, research and teaching. Over the past thirty years, CHEO has established itself as a world-class academic health sciences center providing leading-edge treatment, diagnostic and laboratory services for children and youth aged 0 to 18 years. CHEO houses the Provincial Centre of Excellence for Child and Youth Mental Health and the Ontario Newborn Screening Program. CHEO is an active partner in the Champlain Local Health Integration Network, providing leadership in all aspects of paediatric health and wellbeing. Questions to Mike were uplinked beforehand. “How do you take a shower/bath?”; “What is the best part about being an astronaut?”; “What do you eat?”; “How do you communicate with family?”; “What do you do if you dislike another crew member?”; “What do you do for fun?”; “Do you sleep in a bed? Or do you just “bob”?”; “What kind of clothes do you wear? Is it hot or cold?”; “Have you ever seen a balloon in space?”; “What kind of training do you need before you can live on the ISS?”]

The station residents conducted their regular daily 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (FE-2), TVIS treadmill (CDR, FE-1), ARED/Advanced Resistive Exercise Device (CDR, FE-2) and VELO bike with bungee cord load trainer (FE-1).

The crew also had their periodic PMCs (Private Medical Conferences) via S- & Ku-band audio/video, Yuri at ~9:20am, Mike at ~10:20am, Sandy at ~1:35pm EST.

No CEO (Crew Earth Observations) photo targets uplinked for today.

CEO photography can be studied at this “Gateway” website: (as of 9/1/08, this database contained 770,668 views of the Earth from space, with 324,812 from the ISS alone).

ISS Orbit (as of this morning [1/20/2009], 7:14am EST [= epoch]):
Mean altitude — 357.5 km
Apogee height — 362.7 km
Perigee height — 352.3 km
Period — 91.69 min.
Inclination (to Equator) — 51.64 deg
Eccentricity — 0.00062480
Solar Beta Angle — 16.1 deg (magnitude decreasing)
Orbits per 24-hr. day — 15.72
Mean altitude loss in the last 24 hours — 27 m
Revolutions since FGB/Zarya launch (Nov. 98) — 58272

Significant Events Ahead (all dates Eastern Time, some changes possible!):
02/04/09 — ISS reboost 2
02/09/09 — Progress M-01M/31P undocking & deorbit
02/10/09 — Progress 32P launch
02/12/09 — Progress 32P docking
02/12/09 — STS-119/Discovery/15A launch – S6 truss segment
02/14/09 — STS-119/Discovery/15A docking
02/24/09 — STS-119/Discovery/15A undocking
02/26/09 — STS-119/Discovery/15A landing (nominal)
03/25/09 — Soyuz TMA-14/18S launch
03/27/09 — Soyuz TMA-14/18S docking (DC1)
04/05/09 — Soyuz TMA-13/17S undocking
04/07/09 — Progress 32P undocking & deorbit
05/12/09 — STS-125/Atlantis Hubble Space Telescope Service Mission 4 (SM4)
05/15/09 — STS-127/Endeavour/2J/A launch – JEM EF, ELM-ES, ICC-VLD
05/27/09 — Soyuz TMA-15/19S launch

Six-person crew on ISS
08/06/09 — STS-128/Discovery/17A – MPLM (P), LMC, last crew rotation
08/XX/09 — Soyuz 5R/MRM2 (Russian Mini Research Module, MIM2) on Soyuz
09/XX/09 — H-IIB (JAXA HTV-1)
11/12/09 — STS-129/Atlantis/ULF3 – ELC1, ELC2
12/10/09 — STS-130/Endeavour/20A – Node-3 + Cupola
02/11/10 — STS-131/Atlantis/19A – MPLM(P), LMC
04/08/10 — STS-132/Discovery/ULF4 – ICC-VLD, MRM1
05/31/10 — STS-133/Endeavour/ULF5 – ELC3, ELC4
12/XX/11– Proton 3R/MLM w/ERA.